If the price elasticity of demand for iPhones were 1.8, by how much would iPhone unit sales have increased if the launch price had been $399 instead of $499?
1
Expert's answer
2011-10-06T09:18:12-0400
Ε(Q)=(ΔQ/Q1)/(ΔP/P1)
1,8 = ((Q2-Q1)/Q1)/(499-399)/399)

(Q2-Q1)/Q1 =180/399

Q2/Q1 – 1 =180/399

Q2/Q1 =579/399 = 1,45

iPhone unit sales have increased in 1,45 times.